OpenStudy (anonymous):

i need help:( Write and equation of the liine that passes through point P and parallel to the line with the given equation. P(3,3) Y=-1/2x + 3.....

OpenStudy (aripotta):

ok, so parallel lines have what in common?

OpenStudy (anonymous):

the same slope?

OpenStudy (aripotta):

yes. so we can put the information given to us in point-slope form first, which is: y - y1 = m(x - x1) m would be -1/2, since both lines are parallel. so what does our point-slope form look like after you substitute?

OpenStudy (anonymous):

i dont get it, wait i thought you just PLUG IN THE X AND Y INTO THE EQUATION?

OpenStudy (aripotta):

no, because the two lines don't have the same y-intercept. if you put (3,3) with a slope of -1/2 into point-slope, then make it into slope-intercept, you'll get the right y-intercept

OpenStudy (anonymous):

ok then. but point intercept form is y=mx+b... WHAT DO I DO NOW. Dx

OpenStudy (aripotta):

point-intercept is y - y1 = m(x - x1). slope-intercept is y = mx + b

OpenStudy (anonymous):

i never learned that...

OpenStudy (aripotta):

oh. well, it is :l and standard form is Ax + By = C

OpenStudy (aripotta):

but we're not doing standard form :/

OpenStudy (anonymous):

will you please just tell me what to do :/ ive been at this for like 2 hours now.

OpenStudy (aripotta):